OpenStudy (anonymous):

Find an equation of the lines passing thru each pair of points. Write eqution in form Ax+Bx=C... Here are the points (2,3) and (-1,-1)

OpenStudy (bahrom7893):

Do the same thing we did come on! M = ( -1-3 ) / (-1-2) = 4/3 and then do what we did earlier (y-y1 = M(x-x1), simplify etc..)